OpenStudy (anonymous):

solve the following equation for y -40x+5y=-35

jimthompson5910 (jim_thompson5910):

There are no equal signs. So this isn't an equation

OpenStudy (anonymous):

i ment to put equal not add

OpenStudy (anonymous):

rewrite to solve for y: 5y = 40x +35 y = 8x + 7

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

-40x+5y=-35 5y=40x-35 y=(40x-35)/5 y=(40x)/5-(35)/5 y=8x - 7